# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

### v1

@@ III-C First-level Graph
III-D Second-level Graph
In this section, we construct a factor graph to optimize the variables, ultimately addressing the MAP inference problem introduced in Eq. (3). Suppose that the $j$-th marker is detected in the $i$-th scan, the following nodes are added to the second graph. The variable nodes include $\mathbf{T}^{j}$ (6-DOF pose of the $j$-th marker w.r.t. $\{G\}$), $\mathbf{T}_{i}$ (6-DOF pose of the $i$-th scan w.r.t. $\{G\}$), and $\{\mathbf{p}^{j,1},\mathbf{p}^{j,2},\mathbf{p}^{j,3},\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{G\}$). The factor nodes include $\mathbf{T}^{j}_{i}$ (6-DOF pose of the $j$-th marker w.r.t. $\{F\}_{i}$), $\{^{j}\mathbf{p}^{j,1},\ ^{j}\mathbf{p}^{j,2},\ ^{j}\mathbf{p}^{j,3},\ ^{j}\mathbf{p}^{j,4}\}$ (3D coordinates of the coroners of the $j$-th marker w.r.t. $\{M\}_{j}$), and $\{_{i}\mathbf{p}^{j,1},\ _{i}\mathbf{p}^{j,2},$ ${}_{i}\mathbf{p}^{j,3},$ ${}_{i}\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{F\}_{i}$).
For clarity, the first stage in Fig. 5 shows the added nodes and edges when a marker is detected in a scan. The graph becomes the one depicted in the second stage of Fig. 5 after integrating all the marker detection results. Subsequently, we add a prior factor that connects to the first (anchor) point cloud pose node since we set $\{F\}_{1}$ as $\{G\}$. Finally, we add factor nodes representing the relative poses between the anchor scan and each non-anchor scan. Up to this point, the factor graph is completed, as shown in stage 3 of Fig. 5. Following [30, 31], this indicates that the factorization of the posterior probability $P(\Theta\mid\mathcal{Z})$ has been defined:
$P(\Theta\mid\mathcal{Z})=\prod_{k}P^{(k)}(\Theta),$ (6)
Figure 6: An illustration of $\lambda^{*}$ across various scenes and positions.
We utilize the Levenberg-Marquardt algorithm [31] to resolve this problem. The acquisition of initial values is introduced in Section III-C. The noise covariance matrices are determined by the quantitative experiments conducted in [20].
Remark. To clarify, the difference between IFM [20] and this work resembles the difference between AprilTag [11] and marker-based SfM [15, 13]. Although IFM and AprilTag address the problem of sensor pose estimation, constructing an efficient framework for multiview point cloud registration or SfM based on markers is a different endeavor, which is the contribution of [15, 13] as well as this work.
IV Experimental Results
In this section, we report the experimental results of five tests to demonstrate the effectiveness of the proposed method from different perspectives.
IV-A Evaluation of the Adaptive Threshold Method
Data. We place the Livox MID-40 LiDAR and a 69.2 cm $\times$ 69.2 cm ArUco marker in three different scenes as shown in Fig. 6(a). The three scenes are between groups of buildings, open outdoors, and large indoor parking lots, respectively. In each scene, we collect point clouds and test the proposed adaptive threshold method at different relative positions. Results and Analysis. The 3D stem plots in Fig. 7 present the results, where x and y indicate the relative position of the LiDAR in the marker coordinate system, and the z-axis corresponds to $\lambda^{*}$. In scene 1, we test at 21 different locations, and $\lambda^{*}$ varies from 7 to 36. In scene 2, we test at 21 viewpoints, and $\lambda^{*}$ varies from 6 to 15. In scene 3, we test at 9 positions, and $\lambda^{*}$ varies from 6 to 20. The experiments carried out in this section demonstrate that when the viewpoint changes dramatically, a constant threshold [20] is not applicable. Thus, the proposed adaptive threshold method is necessary.
IV-B Evaluation of Point Cloud Registration Accuracy
Data. Given that the existing point cloud registration benchmark [6] lacks fiducial markers in the scenes, we construct a new test dataset, as shown in Fig. 7, with the Livox MID-40. As illustrated in the caption of Fig. 7, our test datasets cover various indoor and outdoor scenes. Indoors, multiple 16.4 cm × 16.4 cm AprilTags [11] are positioned in the environment. Outdoors, multiple 69.2 cm × 69.2 cm ArUcos [12] are placed in the environment. Note that since the LiDAR fiducial markers in this work are thin sheets of objects attached to other surfaces, they are almost invisible in the point clouds. This is infeasible if we adopt LiDARTags [19] or calibration boards [16, 17, 18] to provide fiducials.
TABLE II: comparison with LOAM methods regarding reconstruction
@@ IV-B Evaluation of Point Cloud Registration Accuracy
Livox Mapping [23] 0.0106 75.27
LOAM Livox [5] 0.0335 78.82
Traj LO [22] 0.0107 82.88
Ours 0.0030 96.22
Competitors and Metrics. Competitors include the state-of-the-art (SOTA) learning-based multiview point cloud registration method, SGHR [1], and the SOTA pairwise geometry-based method, Teaser++ [3]. We employ the root-mean-square errors (RMSEs) [32] as the metric. We also compare the inference time of the methods with an AMD Ryzen 7 5800X CPU. Results and Analysis. The qualitative and quantitative results are presented in Fig. 7 and Table I, respectively. As seen, the proposed method yields the best accuracy and efficiency. SGHR [1] successfully registers the point clouds in scenes 1 to 3 while failing in scenes 4 to 10. Its efficiency is poor on the CPU, while the inference time is around 10s to 20s on an RTX 3090 GPU. Considering that Teaser++ [3] is a pairwise method, we manually select pairs of point clouds with overlap as the inputs (the same scheme adopted in the following tests), yet it still fails in all scenes. The failure cases show SGHR’s limited generalization to unseen scenarios, and Teaser++ struggles with low overlap cases.
Figure 9: Visual comparison of the instance reconstruction. From top to bottom: ground truth, Ours, Traj LO [22], Livox Mapping [23], and LOAM Livox [5].
IV-C Evaluation of Instance Reconstruction Quality
An important application of point cloud registration is the 3D reconstruction for asset collection [33]. Thus, we evaluate instance reconstruction quality in this test. Data. The experimental setup is depicted in Fig. 8(a). Four 69.2 cm $\times$ 69.2 cm ArUco [12] markers are placed in the environment. As depicted by the yellow trajectory, the Livox MID-40 LiDAR follows a looping path to scan the vehicle, pausing at five viewpoints for a few seconds to collect relatively dense point clouds due to its sparse scanning pattern [5]. We record the rostopic of the point cloud published by the LiDAR sensor as a rosbag throughout the sampling process. Then, the same rosbag is provided to all competitors. Point cloud registration methods [1, 3], including our method, only utilize a portion of the rosbag, i.e., the input point clouds shown in Fig. 8.
TABLE III: comparison with competitors regarding localization accuracy in a degraded scene.
[…]
Traj LO [22] 0.8289 0.6077
Livox Mapping [23] 0.4854 0.0893
LOAM Livox [5] 0.8503 0.5921
@@ IV-C Evaluation of Instance Reconstruction Quality
To compare the instance reconstruction quality, we use supervisely [34], a popular 3D annotation tool, to extract the point cloud of the vehicle from each scene reconstruction result. Since the manufacture and model (Mercedes-Benz GLB) of the vehicle to be reconstructed are known, a high-fidelity 3D model acquired from a 3D assets website acts as the ground truth shape. Competitors and Metrics. The competitors are point cloud registration methods (SGHR [1] and Teaser++ [3]) and LOAM methods (Traj LO[22], Livox Mapping [23], and LOAM Livox [5]). Following [35], we compute the Chamfer Distance and recall between the ground truth shape and each reconstruction result to perform a quantitative comparison of the reconstruction quality. Results and Analysis. The qualitative results for point cloud registration methods and LOAM methods are shown in Fig. 8 and Fig. 9, respectively. The quantitative results are presented in Table II. As seen in Fig. 8, SGHR is only able to align the third and fifth scans, while Teaser++ fails in all scenes even when pairs with overlap are manually selected and provided. In terms of instance reconstruction quality, as depicted in Fig. 9, our result preserves intricate details. The result of Traj LO [22] has almost no drift while the surface is noisy. The Livox Mapping [23] result exhibits noticeable drift and surface noise, while LOAM Livox [5] demonstrates severe drift. As shown in Table II, our method achieves the highest reconstruction quality.
TABLE IV: Comparison with KISS-ICP [4] regarding localization.
@@
Figure 10: Visualization of the scene reconstruction results of different methods.
IV-D Evaluation in a Degraded Scene
As previously introduced, the adoption of fiducial markers can enhance the robustness of reconstruction and localization in degraded scenes. Therefore, in this test, we evaluate our method in a degraded scene. Data. The setup is shown in the reference of Fig. 10. This scenario has repetitive structures and weak geometric features. We attach thirteen 16.4 cm $\times$ 16.4 cm AprilTags to the wall. The LiDAR follows an arc trajectory to scan the scene, pausing at 11 viewpoints for a few seconds to collect 11 dense point clouds. We record the point cloud rostopic as a rosbag and provide it to all competitors, excluding SfM-M [15]. Point cloud registration methods only leverage a portion of the rosbag, consisting of the 11 point clouds shown in the input of Fig. 10. We also captured 72 images with an iPhone 13 to use as input for SfM-M. The ground truth trajectories are given by an OptiTrack Motion Capture system. Competitors and Metrics. Considering that all point-cloud-based competitors [1, 3, 22, 23, 5] struggle in the degraded scene and there does not exist a point-cloud-based competitor that can utilize our fiducial markers (thin sheets of letter-size paper), we add SfM-M [15], the state-of-the-art marker-based SfM method, as a competitor. This addition enables a comprehensive and meaningful comparison. We employ the RMSEs as the metric. Results and Analysis. The qualitative results and trajectory comparison are presented in Fig. 10. As depicted in Fig. 10, only our method and SfM-M [15] successfully reconstruct the scene, while the other competitors [1, 3, 22, 23, 5] fail in this degraded scene. The quantitative results shown in Table III demonstrate that the proposed approach achieves the best localization accuracy. Moreover, the performance of SfM-M is inferior to our method, which is reasonable considering LiDAR is a ranging sensor.
IV-E Evaluation of Outdoor Localization
Localization is also a crucial application of point cloud registration methods. Thus, we compare the proposed method against KISS-ICP [4] in terms of outdoor localization in this test. As shown in Fig. 11(a), Four 69.2 cm $\times$ 69.2 cm ArUco [12] markers are deployed in the environment. The vehicle, equipped with an RS-Ruby 128-beam mechanical LiDAR, follows an 8-shaped trajectory without pausing and samples 364 LiDAR scans. The ground truth is obtained from Real-Time Kinematic. As seen in Fig. 11(b) and Table IV, KISS-ICP and our method show comparable performance, with our method exhibiting less drift in the middle of the trajectory.
V Conclusion
@@
In this work, we develop L-PR, a novel framework that exploits LiDAR fiducial markers for unordered low overlap multiview point cloud registration. L-PR consists of an adaptive threshold marker detection method and two levels of graphs. The markers in this work are thin sheets of paper attached to other surfaces, without impacting the 3D environment. The adaptive threshold method robustly detects markers even when there are dramatic changes in viewpoints among scans, thus overcoming the limitation of [20]. The proposed two-level graph efficiently aligns the low overlap multiview point clouds directly from an unordered set by optimizing a MAP problem. We conduct qualitative and quantitative experiments to demonstrate the superiorities of L-PR over competitors in terms of (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene.
Limitations and Discussions. While our proposed method shows promising results, it is constrained to scenes with deployed fiducial markers, a limitation shared by all marker-based approaches [15, 13]. Nevertheless, marker-based methods typically prioritize controllable environments, aiming to enhance reconstruction and localization in scenarios with low overlap and/or lack of texture. From this perspective, our method, developed for LiDARs, achieves the same core objective as previous methods [15, 13] designed for cameras. Moreover, as depicted in Fig. 7, the adoption of markers has no effect on the 3D geometry of the reconstructed result, given that the markers used in this work are thin sheets of paper attached to other planes.

### v2

@@ V-B Evaluation of Point Cloud Registration Accuracy
TABLE III: quantitative evaluation of the enhancement of SGHR [1] due to the addition of Livox-3DMatch to the training data.
Method Dataset RR $\uparrow$ Dataset RR $\uparrow$ Dataset $\mathrm{RMSE}_{T}$ (m) $\downarrow$ $\mathrm{RMSE}_{R}$ (deg) $\downarrow$
(%) (%) Mean/Med Mean/Med
SGHR [1] 3DMatch [6] 92.68 ETH [21] 93.74 ScanNet [22] 0.66/0.51 23.50/22.08
SGHR [1] + Livox-3DMatch (our data) 95.58 98.03 0.51/0.45 20.87/17.21
V-C Application 1: 3D Asset Collection
An important application of point cloud registration is the 3D asset collection [34]. Thus, we evaluate the instance reconstruction quality of L-PR in this test. Data. The experimental setup is depicted in Fig. 9. Four 69.2 cm $\times$ 69.2 cm ArUco [11] markers are placed in the environment. We employ the Livox MID-40 LiDAR to scan the vehicle from five significantly different viewpoints. As shown in Fig. 9, this is also a challenging low-overlap case, with the average and minimum overlap rates being 13.39% and 1.02%, respectively. To evaluate the instance reconstruction quality, we use Supervisely [35], a popular 3D annotation tool, to extract the point cloud of the vehicle from the reconstruction result. Since the manufacture and model (Mercedes-Benz GLB) of the vehicle to be reconstructed are known, a high-fidelity 3D model acquired from a 3D assets website acts as the ground truth shape. Metrics. Following [34], we compute the Chamfer Distance and Recall between the ground truth shape and the reconstruction result. Results and Analysis. The qualitative result is shown in Fig. 9. As seen, SGHR is only able to align the third and fifth scans, while Teaser++ fails in all scenes even when pairs with overlap are manually selected and provided. By contrast, L-PR successfully aligns the unordered multiview point clouds and our reconstructed result preserves intricate details. The Chamfer Distance and Recall of our reconstructed result are 0.003 and 96.22%, respectively. This indicates [34] that L-PR can serve as a convenient tool for collecting high-fidelity 3D assets using the LiDAR sensor.
V-D Application 2: Training Data Collection and Enhancement of Existing Learning-Based Method
Training data is crucial for learning-based methods. Most existing methods [1, 2] are trained on the 3DMatch dataset [6]. Augmenting the training data is beneficial for improving the generalization ability of learning-based methods. As demonstrated in Sections V-B and V-C, L-PR can automatically and efficiently register multiview low overlap 3D point clouds. Thus, it can serve as a valuable tool for collecting training data. In this section, we demonstrate that the training data collected using L-PR can enhance the performance of the SOTA method, SGHR [1], on various benchmarks, including 3DMatch [6], ETH [21], and ScanNet [22]. In particular, using L-PR, we process all the point clouds shown in Figs. 8 and 9, including 11 scenes with 33 scans, into the format required for training by SGHR. The newly collected data is named Livox-3DMatch. Then, we train SGHR from scratch using a mix of 3DMatch and Livox-3DMatch. In particular, the number of pairs in the training data increased from 14,400 to 17,700, an increase of 22.91%, thanks to the introduction of our data. Finally, we compare the performance of the SGHR models trained with only 3DMatch and with 3DMatch plus our data on popular benchmarks [6, 21, 22]. Metrics. Following [1], we use Registration Recall (RR) to evaluate performance on 3DMatch [6] and ETH [21], and RMSEs to report performance on ScanNet [22]. Results and Analysis. The quantitative results are presented in Table III. As seen, the RR of SGHR is increased by 2.90% on 3DMatch [6] and 4.29% on ETH [21]. The translation error and rotation error on ScanNet [22] are decreased by 22.72% and 11.19%, respectively.
V-E Application 3: Reconstructing a Degraded Scene
As previously introduced, the use of fiducial markers can enhance the robustness of reconstruction and localization in degraded scenes. Therefore, in this test, we evaluate our method in a textureless, degraded scene. Data. The setup is shown in the reference of Fig. 10. This scenario has repetitive structures and weak geometric features. We attach thirteen 16.4 cm $\times$ 16.4 cm AprilTags to the wall. The LiDAR scans the scene from 11 viewpoints. We also captured 72 images with an iPhone 13 to use as input for SfM-M [14]. The ground truth trajectories are given by an OptiTrack Motion Capture system. Competitors and Metrics. Considering the competitors [1, 3] struggle in the degraded scene, we add SfM-M [14], the state-of-the-art marker-based SfM method, as a competitor. This addition enables a comprehensive and meaningful comparison. We employ the RMSEs as the metric. Results and Analysis. The qualitative results and trajectory comparison are presented in Fig. 10. Once again, SGHR and Teaser++ struggle in this scene. The quantitative results shown in Table IV demonstrate that the proposed approach achieves better localization accuracy, which is expected given that LiDAR is a ranging sensor.
Figure 10: Visualization of the scene reconstruction and localization results of different methods.
TABLE IV: comparison with SfM-M [14] regarding localization accuracy in a degraded scene.
@@ V-E Application 3: Reconstructing a Degraded Scene
V-F Application 4: Localization in a GPS-denied Environment
Localization is also a crucial application of point cloud registration methods. Fiducial markers are a popular tool for providing localization information in GPS-denied environments, such as indoor parking lots. In this test, we compare the proposed method with KISS-ICP [4] in terms of localization accuracy. As shown in Fig. 11(a), four 69.2 cm $\times$ 69.2 cm ArUco [11] markers are deployed in the environment. The vehicle, equipped with an RS-Ruby 128-beam mechanical LiDAR, follows an 8-shaped trajectory without pausing and samples 364 LiDAR scans. We conduct the experiment on the roof of a large parking lot to acquire the ground truth trajectory from the Real-Time Kinematic. However, our method is also applicable to indoor GPS-denied environments. As seen in Fig. 11(b) and Table V, our method exhibits less drift in the middle of the trajectory.
TABLE V: Comparison with KISS-ICP [4] regarding localization.
@@ V-F Application 4: Localization in a GPS-denied Environment
V-G Application 5: 3D Map Merging
By cooperating with the algorithm introduced in [36], the proposed L-PR can be utilized for 3D map merging, as shown in Fig. 12. These 3D maps are extremely low overlap cases. In particular, the overlap rates [2] are 4.87% between map 1 and map 2, 3.96% between map 1 and map 3, and 2.59% between map 2 and map 3.
Figure 12: An illustration of 3D map merging using L-PR.
VI Conclusion
@@
In this work, we develop L-PR, a novel framework that exploits LiDAR fiducial markers for unordered low overlap multiview point cloud registration. The markers in this work are thin sheets of paper attached to other surfaces, without impacting the 3D environment. L-PR consists of an adaptive threshold marker detection method and two levels of graphs. The adaptive threshold method robustly detects markers even when there are dramatic changes in viewpoints among scans, thus overcoming the limitation of [19]. The proposed two-level graph efficiently aligns the low overlap multiview point clouds directly from an unordered set by optimizing a MAP problem. We conduct qualitative and quantitative experiments to demonstrate the superiority of L-PR over previous SOTA methods and the diverse applications of L-PR, including 3D asset collection, training data collection, reconstruction of a degraded scene, localization in a GPS-denied environment, and 3D map merging. In particular, we collect a new training dataset named Livox-3DMatch using L-PR. The Livox-3DMatch dataset expands the original 3DMatch training data from 14,400 pairs to 17,700 pairs, representing a 22.91% increase. Training on this augmented dataset improves the performance of the SOTA learning-based method SGHR [1] by 2.90% on 3DMatch [6], 4.29% on ETH [21], and 22.72% (translation) / 11.19% (rotation) on ScanNet [22]. Future work includes utilizing L-PR to collect more 3D assets and point cloud registration training data to benefit the development of learning-based methods [34, 1].